# Heterogeneous Graph Neural Networks for End-to-End Traffic Assignment and Traffic Flow Learning

### v2

@@ 4.2 Experiments on Generalized Synthetic Networks
In this section, unlike the experiments in Section 4.1, which involved training and testing on an identical network topology, we examine the generalization capability of the proposed model to networks with varied topologies. In particular, we consider real-world scenarios in which certain links in the network are fully closed due to governmental directives or catastrophic events such as bridge collapses, leading to significant alterations in network topology. Another scenario for topology alternation is when cities consider network expansion to better serve increased mobility demands due to current urbanization trends. Under all these scenarios, the resulting urban networks may exhibit both commonalities and disparities in their topologies (Rodrigue, 2020). This is while training models separately for each distinct network requires substantial time and effort. Motivated by the aforementioned considerations, our aim is to explore the generalization ability of our HetGAT model over varying topologies.
@@
In this work, we conduct experiments on two sets of networks: (1) modified urban networks of Section 4.1; (2) synthetic networks. The first set, includes networks that are modifications of the original Sioux Falls, EMA, and Anaheim networks by adding and removing links. For each of these networks, we generate 20 unique topological variations. Additionally, we incorporate three configurations of OD demand in our experiments: complete OD setting, 20% and 40% incomplete OD setting. The training setting and the hyperparameters remain the same as those in the aforementioned experiments in Section 4.1. The results are presented in Table 4. It can be seen that compared to the results in Table 2, HetGAT exhibits comparable levels of effectiveness for all networks. Additionally, in comparison to other baseline models, our proposed HetGAT model consistently outperforms both GAT and GCN over all networks, under both complete and incomplete OD scenarios. It is noteworthy that compared to the homogeneous GNN model, since our model has captured the influence of OD pairs via virtual links, the HetGAT model is more expressive and can better capture and learn the impact of topology alterations on flow distributions.
@@
The second set includes synthetic networks. We generate these networks by starting with a grid graph and then adding links between randomly selected nodes. Furthermore, to emulate the real road network, a number of nodes and edges are randomly removed until the number of nodes in the graph reaches a predefined threshold. Two sets of synthetic networks are considered: one with 100 nodes and another with 300 nodes. For each graph size, 20 different graph topologies are generated. Three of these randomly generated graphs of size 100 are shown in Figure 8. The OD demand and the link capacity are also randomly generated using the scaling factor according to Equations 15 and 16, respectively. To demonstrate the diversity of these examples, similar to Section 4.1, the histograms of the link capacities and the OD demands in the generated dataset are shown in Figure 7, indicating the training and testing data are sufficiently diverse to cover different scenarios.
In this work, we consider a case in which the training and test graphs are of the same size, and a case where the training and test graphs are of different size (e.g., training graphs of size 100 and test graphs of size 300). It should be noted that the proposed HetGAT involves a node feature whose dimension is equal to the total number of nodes in the graph. Therefore, this requires an adjustment when training and testing are done on graphs with different number of nodes. In additional to standard training, which trains and tests on the same set of graphs, we propose two training strategies, enabling the HetGAT model to be trained and tested effectively on various graphs:
@@
    Transfer learning: we adapt the model trained from previous experiments by substituting the preprocessing and final layers with new parameters, while freezing the remaining parameters unchanged. This facilitates a quick adaption to the new topologies with limited training.
@@
    Homogenized training: we define a maximum node number threshold $N_{\max}$ and for all the graphs graphs containing fewer nodes than this threshold, we introduce dummy nodes to the graph until the number of node reaches the $N_{\max}$. This approach ensures uniform formulation among different graphs, allowing a single model to train and test across graphs of varying sizes.
The training setting and the hyperparameters remain the same as those in the previous experiments. The prediction performance metrics for the testing sets are presented in Table 5. Compared with transfer learning and homogenized learning settings, the proposed model achieves the highest accuracy when same graph size is used in training and testing. Despite a performance decline under the transfer and homogenized learning setting, the HetGAT model maintains a competitive edge, outperforming baseline models such as GAT and GCN.
Table 5: Comparison of the performance of HetGAT with that of GAT and GCN on generalized synthetic networks. Two graph sizes of 100 and 300 are considered in the experiments. We consider three different training strategies: standard training, transfer learning, and homogenized learning.
Network Model Standard training Transfer learning Homogenized learning
@@
300 GCN 2.97 4.11 31.87% 44.14% 0.15 3.09 4.28 33.17% 45.97% 0.13 3.04 4.23 32.65% 45.32% 0.14
@@ 5 Conclusion and Discussion
In this paper, we proposed a novel approach for end-to-end traffic assignment and traffic flow learning using heterogeneous graph neural networks. Compared with conventional homogeneous graph neural networks, we introduce the auxiliary virtual links connecting origin-destination node pairs. Furthermore, we proposed a novel adaptive graph attention mechanism to effectively capture the semantic and contextual features through different types of links. Furthermore, we conducted extensive experiments on three real-world urban transportation networks to evaluate the performance of the proposed model and compared it with the state-of-the-art models. The results indicate that the proposed model outperforms other models in terms of convergence rate and prediction accuracy. Notably, by introducing two different training strategies, the proposed heterogeneous graph neural network model can also be generalized to different network topologies, underscoring its potential in real-world scenarios.
Our work can be considered as an efficient surrogate model for further accelerating the complex optimization tasks including resource allocation and infrastructure asset management. Currently, the proposed HetGAT model only learns and predicts the static traffic flow patterns. As a potential extension of this work, the proposed framework can be extended to learn the dynamic traffic flow patterns. Furthermore, the current proposed GNN model uses training data collected from conventional solvers of static traffic assignment. In future work, we will explore how these models can be trained on traffic data collected from sensors, such as loop detectors, cameras, or GPS devices.
Acknowledgment

### v3

@@ 5.2 Experiments on Generalized Synthetic Networks
In this section, unlike the experiments in Section 5.1, which involved training and testing on an identical network topology, we examine the generalization capability of the proposed model to networks with varied topologies. In particular, we consider real-world scenarios in which certain links in the network are fully closed due to governmental directives or catastrophic events such as bridge collapses, leading to significant alterations in network topology. Another scenario for topology alternation is when cities consider network expansion to better serve increased mobility demands due to current urbanization trends. Under all these scenarios, the resulting urban networks may exhibit both commonalities and disparities in their topologies (Rodrigue, 2020). This is while training models separately for each distinct network requires substantial time and effort. Motivated by the aforementioned considerations, our aim is to explore the generalization ability of our HetGAT model over varying topologies. In this work, we conduct experiments on two sets of networks: (1) modified urban networks of Section 5.1; (2) synthetic networks. The first set, includes networks that are modifications of the original Sioux Falls, EMA, and Anaheim networks by adding and removing links.
5.2.1 Modified urban networks
In this section, we generate synthetic networks by modifying the three urban networks used in previous seciton. For each of these networks, we generate 20 unique topological variations. Additionally, we incorporate three configurations of OD demand in our experiments: complete OD setting, 20% and 40% incomplete OD setting. The training setting and the hyperparameters remain the same as those in the aforementioned experiments in Section 5.1. The results are presented in Table 4. It can be seen that compared to the results in Table 2, HetGAT exhibits comparable levels of effectiveness for all networks. Additionally, in comparison to other baseline models, our proposed HetGAT model consistently outperforms both GAT and GCN over all networks, under both complete and incomplete OD scenarios. It is noteworthy that compared to the homogeneous GNN model, since our model has captured the influence of OD pairs via virtual links, the HetGAT model is more expressive and can better capture and learn the impact of topology alterations on flow distributions.
@@ 5.2.1 Modified urban networks
5.2.2 Randomly generated graphs and generalization to variable graph size
The second set of synthetic networks are randomly generated by starting with a grid graph and then adding links between randomly selected nodes. Furthermore, to emulate the real road network, a number of nodes and edges are randomly removed until the number of nodes in the graph reaches a predefined threshold. Two sets of synthetic networks are considered: one with 100 nodes and another with 300 nodes. For each graph size, 20 different graph topologies are generated. Three of these randomly generated graphs of size 100 are shown in Figure 7. The OD demand and the link capacity are also randomly generated using the scaling factor according to Equations 15 and 16, respectively. To demonstrate the diversity of these examples, similar to Section 5.1, the histograms of the link capacities and the OD demands in the generated dataset are shown in Figure 8, indicating the training and testing data are sufficiently diverse to cover different scenarios.
@@ 5.2.2 Randomly generated graphs and generalization to variable graph size
In this section, we investigate the generalization of the proposed method when graph sizes in the training and test cases are different. In particular, we consider two different graph sizes, with the node numbers $N=100$ and $N=300$. It should be noted that a node feature in our proposed HetGAT model has a dimension equal to the total number of nodes in the graph. Therefore, to handle variable graph sizes in the dataset, we need to make an adjustment to the model. To this end, we propose two training strategies for HetGAT:
@@
    Transfer learning: In this approach, we use the model trained from previous experiments as a pre-trained model and let the parameters of its preprocessing and final layers be re-trained. The remaining parameters will be frozen and unchanged during this re-training. So, for subsequent training involving graphs of the new size, the optimization will be relatively fast as it is only done for the parameters in two layers, rather than the entire model. Two cases are studied: when testing is done on graphs with $N=100$ and training is done with $N=300$, and vice versa.
@@
    Homogenized training: In this approach, we set a ceiling (maximum) for the node number $N_{\max}$, anticipating the largest graph that can ever be handled. Then, for all the graphs containing fewer nodes, we add enough dummy nodes to the graph for it size to reach $N_{\max}$. This approach ensures uniform formulation and architecture across different graphs, allowing a single model to be used for graphs of varying sizes. To numerically study this strategy, the model is trained once using graphs with sizes $N=100$ and $N=300$. This trained model is then tested on graphs with $N=100$ and graphs with $N=300$ separately.
In order to evaluate the effectiveness of our proposed model compared to baseline models, we compare our results in a standard training, where training and test graphs are of the same size. That means for test graphs with $N=100$, HetGAT was trained in the standard way of the previous section, with only graphs with $N=100$. Similarly, training and test graphs of size $N=300$ were used for the second set of “standard training” results.
The training setting and the hyperparameters remain the same as those in the previous experiments. The prediction performance metrics for the testing sets are presented in Table 5 and 6. First, it can be seen that the proposed HetGAT model achieves the best accuracy levels compared to other model architectures on randomly generated graphs when same graph size is used in training and test. Furthermore, for the cases with variable graph sizes, even though the performance decline under the transfer and homogenized learning setting, the proposed HetGAT model maintains a competitive edge, outperforming baseline models such as GAT and GCN.
Table 5: Comparison of the performance of HetGAT with that of GAT and GCN on generalized synthetic networks. The graph size in testing dataset is 100. We consider three different training strategies: standard training, transfer learning, and homogenized learning.
Testing dataset Model Standard training Transfer learning Homogenized learning
Training N = 100 Training N = 300 Training N = {100, 300}
@@
Table 6: Comparison of the performance of HetGAT with that of GAT and GCN on generalized synthetic networks. The graph size in testing dataset is 300. We consider three different training strategies: standard training, transfer learning, and homogenized learning.
Testing dataset Model Standard training Transfer learning Homogenized learning
Training N = 300 Training N = 100 Training N = {100, 300}
Flow
($1\times 10^{2}$)
Link
utilization
$\tilde{L}_{c}$
Flow
($1\times 10^{2}$)
Link
utilization
$\tilde{L}_{c}$
Flow
($1\times 10^{2}$)
Link
utilization
$\tilde{L}_{c}$
MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE
N=300 GCN 2.97 4.11 31.87% 44.14% 0.15 3.09 4.28 33.17% 45.97% 0.13 3.04 4.23 32.65% 45.32% 0.14
@@ 6 Conclusion and Discussion
In this paper, we proposed a novel approach for traffic assignment using an end-to-end heterogeneous graph neural network. Compared to conventional homogeneous graph neural networks, our proposed architecture includes additional virtual links connecting origin-destination node pairs, to better uncover dependencies between link flows and OD demand. We also proposed a novel adaptive graph attention mechanism to effectively capture the semantic and contextual features through different types of links. The extensive experiments on three real-world urban transportation networks showed that the proposed model outperforms other state-of-the-arth models in terms of convergence rate and prediction accuracy. Notably, by introducing two different training strategies, the proposed heterogeneous graph neural network model can also be generalized to different network topologies, underscoring its potential in real-world scenarios.
The proposed framework in this work can serve as a surrogate model that can significantly accelerate complex optimization tasks in areas such as resource allocation and infrastructure asset management. The significant reduction in computational time allows planners to analyze networks under a large number of scenarios in a more comprehensive and more realistic decision-making process. The generalization advantage of this model is particularly beneficial for network design, especially considering network expansion options. Moreover, it was demonstrated the HetGAT model can robustly predict link flows from inaccurate OD demand data. In this current version, the proposed HetGAT model only learns and predicts the static traffic flow patterns. As a potential extension of this work, the proposed framework can be extended to learn the dynamic traffic flow patterns. Furthermore, the current proposed GNN model uses training data collected from conventional solvers of static traffic assignment. In future work, we will explore how these models can be trained on traffic data collected from sensors, such as loop detectors, cameras, or GPS devices.
7 Acknowledgment